OpenStudy (anonymous):

(x-6)/(6x-1) Someone wanna help me?

OpenStudy (anonymous):

hey:)

OpenStudy (anonymous):

how are you?

OpenStudy (anonymous):

good.. so what do you need to do ?

OpenStudy (anonymous):

ok but what about it ? nothing much to simplify (1/6) - (35/(36x-6))

OpenStudy (anonymous):

maybe you need domain ?

OpenStudy (anonymous):

yep so the excluded value is given if you divide by zero 6x-1 = 0 find x from here..

OpenStudy (anonymous):

the answer is 1/6 in fact but it is coincidence that it came the same as i wrote up there ..

OpenStudy (anonymous):

yes .. you cant divide by zero .. if x = 1/6 you divide by zero so we exclude 1/6 from our x values

OpenStudy (anonymous):

how you divide two fractions ?

OpenStudy (anonymous):

i mean the division of two fractions is like multiplying by reciprocal : \[\frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

agree ?

OpenStudy (anonymous):

we can eliminate (x+3) BUT BEFORE WE DO IT we have to say what the excluded values of x and then eliminate

OpenStudy (anonymous):

do you understand how i got the last line ? can you tell what the excluded x values ?

OpenStudy (anonymous):

i had 3x+9 so taking 3 out gives 3(x+3)

OpenStudy (anonymous):

12x(x+3) = 0 for what x values does this equation hold?

OpenStudy (anonymous):

no.. by hand

OpenStudy (anonymous):

yes 0, -3

OpenStudy (anonymous):

now you can eliminate (x+3) and get your simplified expression

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

sure , it's enough ?

OpenStudy (anonymous):

wait.. you flip the second one

OpenStudy (anonymous):

ok ?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

bye see you